oliviac1111:

What number should be added to both sides of the equation to complete the square? x2 – 10x = 7

Extrinix:

so let's solve this, \(x^2-10x=7\) so first off we need to get a total value of x, so square \(x^2\) and \(7\) \(\sqrt{7} \approx 2.65\) \(x-10x=2.65\) combine like terms \(-9x=2.65\) divide and youll have your answer.

carmelle:

@extrinix wrote:
so let's solve this, \(x^2-10x=7\) so first off we need to get a total value of x, so square \(x^2\) and \(7\) \(\sqrt{7} \approx 2.65\) \(x-10x=2.65\) combine like terms \(-9x=2.65\) divide and youll have your answer.
That is incorrect, we are completing the square, not solving for x. sorry

carmelle:

To start off, we have to divide the coefficient of the x term (which is 10) by 2. When we do that, we get 5. Next, we square 5. `5 x 5 = 25` Now when we plug 25 in to our equation we have: `x^2 + 10x + 25 = 7` But, we must remember, anything we do on one side of an equation, we must do to the other side. So, we add 25 on the other side. Now we have: `x^2 + 10x + 25 = 7 + 25` Since the question asked `“What number should be added to both sides”`, we can conclude, that `we add 25 to both sides`. Hope that helps!! :)